# Federated Learning: A new frontier in the exploration of multi-institutional medical imaging data

### v1

@@ 2.2 Descriptive statistical analysis
The review integrates applications of FL with a wide range of MI modalities, particularly 53 papers related to MRI, 21 to X-ray imaging, 21 to CT, and 15 articles linked to histopathological imaging. In particular, seven articles examined multimodal imaging techniques that involved CT and MRI, CT and PET, CT and X-ray, ultrasonography and mammography, or CT, MRI, ultrasonography and histopathological images. The remaining articles covered PET, OCT, mammography, dermatoscopic, retinal, endoscopic and microscopy imaging. Fig. 4A. provides an exact distribution of articles involving specific topics. Fig. 4B shows a distribution of the papers examined according to the task they address. Among them, 52 articles focus on classification, 43 on segmentation, and 18 on detection. The rest of the papers broadly address prediction, prognosis, reconstruction, translation, restoration, denoising, screening, and super-resolution, with a few considering multiple tasks. Table 2 presents the comparative and detailed information gathered from the evaluated articles, including tasks, employed datasets with references, the presence of heterogeneous data, and the number of clients involved in the analysis.
@@
IXI dataset: https://brain-development.org/ixi-dataset/
Brain Tumor Detection dataset: https://www.kaggle.com/datasets/navoneel/brain-mri-images-for-brain-tumor-detection
Table 2: Summary of federated learning frameworks used in various medical imaging tasks over different imaging modalities. The column “Heterogeneous data” informs whether the heterogeneous data across were used in the aggregation procedure. The last column refers to the number of clients involved in the deep learning task. All links were verified on March 10, 2025.
Reference Task Used datasets Heterogeneous Number
data of clients
@@
Wu et al. [2023b] Reconstruction fastMRI [Knoll et al., 2020], CC359 [Beauferris et al., 2022], own dataset Yes 3
@@
COVID-19-9: https://figshare.com/authors/MedSeg/9940190
MS COVID-19-CT: https://sirm.org/COVID-19/
@@
Wu et al. [2023a] Classification RSNA ICH Detection [Flanders et al., 2020] Yes 20
@@
Zhou et al. [2023] Denoising Dataset from article [Xue et al., 2021], own dataset Yes 3
@@
Naumova et al. [2024] Classification CheXpert [Irvin et al., 2019] N/A 4
@@
Dong and Voiculescu [2021] Detection CheXpert [Irvin et al., 2019], ChestX-ray8 [Wang et al., 2017], VinDr-CXR [Nguyen et al., 2022c] N/A 3, 6
@@
Table 5: Part IV.
Reference Task Used datasets Heterogeneous Number
data of clients
@@
Wu et al. [2023a] Classification HAM10000 [Tschandl et al., 2018], PH2 [Mendonça et al., 2013], Interactive atlas of dermoscopy, ISIC2019 [Codella et al., 2018, Tschandl et al., 2018, Combalia et al., 2019] Yes 5, 10
@@
Agbley et al. [2021] Detection SIIM-ISIC Melanoma Classification dataset N/A 5
@@
Table 6: Part V.
Reference Task Used datasets Heterogeneous Number
data of clients
@@ 3.1 Theory
where $\mathcal{L}$ denotes the global loss function that is computed by combining $C$ local losses $\{\mathcal{L}_{c}\colon c\in\{1,2,...,C\}\}$, each contributing under normalized weight $w_{c}\geq 0$, and $\mathcal{D}_{c}=\left\{\mathbf{x}_{n},\mathbf{y}_{n}\right\}_{n=1}^{n_{c}}$ is the private dataset available only for $c$–th client with training data $\mathbf{x}_{n}$ and corresponding labels $\mathbf{y}_{n}$.
@@ 3.3 Optimization and aggregation
Assuming a local (single) update for client $c$–th performed during one local epoch is given by $\mathbf{\Theta}_{c}^{r}=\mathbf{\Theta}^{r}-\eta\nabla\mathcal{L}_{c}(\mathbf{\Theta}^{r};\mathcal{D}_{c})$, the Eq. (2) can be rewritten as:
@@
$\min_{\mathbf{\Theta}_{c}^{r}}\,\mathcal{L}_{c}(\mathbf{\Theta}_{c}^{r};\mathcal{D}_{c})+\frac{\mu}{2}||\mathbf{\Theta}_{c}^{r}-\mathbf{\Theta}^{r}||_{2}^{2}$ (4)
@@
where $\mathbf{\Theta}_{c,l}^{r}$ refers to the parameters at a global round $r$ from $l$–th non-normalization layer in client $c$.
@@ 3.4 Learning
Vertical FL, also known as heterogeneous FL, handles different feature spaces from the data (e.g., various medical modalities) but with overlapping identifiers, meaning that the same patient appears in all sets [Mothukuri et al., 2021, Brecko et al., 2022, Abad et al., 2022]. It can facilitate training a common distributed model, such as a risk management model [Brecko et al., 2022]. Implementing vertical FL is more challenging than horizontal FL, as it typically involves three types of participants: active, passive, and coordinator. The active participant is the institution that wants to build the model and provide data with both the sample set and their labels. The passive participant only has the sample set, and its feature set differs from the active participant’s. The coordinator is a trusted third party responsible for coordinating training and communication among the participants. In some cases, the active participant is also the coordinator [Li et al., 2023a]. Furthermore, vertical FL requires encryption and a specific learning protocol, as the training process differs from horizontal FL [Mothukuri et al., 2021, Brecko et al., 2022]. Firstly, it is essential to align the training data in order to identify common sample identifiers. After alignment, the participants begin training the model. The most common training protocol is gradient descent, which involves transmitting intermediate values (e.g., softmax after the forward computation). Passive participants send the encoded intermediate results to the active participant. The active participant then calculates the losses and combines them with its loss. The aggregated loss is sent back to the passive participants. Active and passive participants use this aggregated loss to perform backpropagation and update the gradient for their local models. The encrypted gradients are sent to the coordinator, which decrypts them using a private key and returns them to the individual participants. The participants then retrain the model, and this entire process is iteratively repeated until the convergence of the model or a predefined condition is satisfied [Li et al., 2023a, Liu et al., 2024b]
@@ 4.1 Model accuracy
The bias that favors a specific client is another factor that could affect the accuracy of the global model [Darzidehkalani et al., 2022]. If such a situation occurs, it can lead to an increase in the accuracy of the DL task for one client and a decrease in the accuracy for the others. The reason why the bias appears is the variability in the size and distribution of the datasets across the involved sites. The models generated by participants with larger datasets might be assigned with higher weighting factors, thus resulting in an over-representation or under-representation of specific patterns [Abay et al., 2020]. Besides, depending on the aggregation technique, only some clients can be involved, e.g., the institutions that conduct the entire local training within the specific timeframe. This can lead the FL algorithm to favor clients with a better computing infrastructure or faster and more reliable network connections.
@@
4.2 Data heterogeneity
Data heterogeneity is a factor that significantly affects the accuracy of the global model in MI. Imaging data across institutions are typically collected with the equipment provided by various vendors under different acquisition protocols or even various data reconstruction procedures [Gibson et al., 2018, Jiang et al., 2022, Darzidehkalani et al., 2022]. For example, histological images may look different due to various staining conditions. The MRI data may exhibit variations in distributional properties due to multifarious imaging protocols or contrasts applied, such as $T_{1}$-/$T_{2}$-weighting or diffusion-weighting [Jiang et al., 2022], or even the procedure used to reconstruct magnitude data resulting in various statistical properties of the data [Den Dekker and Sijbers, 2014, Pieciak et al., 2016, 2017, Vegas-Sánchez-Ferrero et al., 2017]. Consequently, the data across the institutions may be nonindependent and nonidentically distributed (non-IID).
@@ 4.2 Data heterogeneity
As mentioned above, data heterogeneity originates from the absence of unified imaging standardization procedures. Although the involved sites use the same acquisition protocol to collect data, they may employ different software to preprocess data, leading, for example, to reduced reproducibility of experiments [Veraart et al., 2023]. In addition to these, the number of samples and their distribution can vary over time [Darzidehkalani et al., 2022, Aouedi et al., 2022]. For example, the number of patients affected by a particular disease in one hospital can fluctuate, including the patient’s recovery, death, or transfer of care. Finally, the institutional data deviates according to the demography of the sample. As an illustration, the MRI data change to such an extent throughout life that it allows reconstruction of the patient’s age [Cox et al., 2016, Peng et al., 2021, Pieciak et al., 2023]. Finally, the diverse nature of the data can result in reduced global model accuracy due to its inability to converge. The global model may perform exceptionally well for some clients but inaccurately for others [Darzidehkalani et al., 2022]. According to Zhao et al. [2018], highly heterogeneous data can cause up to 50% degradation in the accuracy of the FL global model, making heterogeneous data one of the most critical issues to address.
In addition to this, institutions still lack proper infrastructure for image data processing and data management systems that are compliant with current standards [Darzidehkalani et al., 2022].
@@ 4.3 Malicious clients
The quality deterioration of the results generated from the global model may come from malicious clients in all institutions. Due to its distributed nature, the FL is susceptible to model poisoning attacks. These attacks occur if a malicious client sends manipulated model updates to the server [Alkhunaizi et al., 2022, Zhang et al., 2022c, Onsu et al., 2023]. Model poisoning attacks can be classified into two categories: untargeted and targeted. In the former case, a corrupted global model generates incorrect predictions for multiple test inputs. In the latter scenario, a corrupted global model causes incorrect predictions for specific test inputs chosen by the attacker, while the accuracy of the global model for other inputs generally remains unaffected [Abad et al., 2022, Zhang et al., 2022c]. For MI applications, the risks posed by malicious clients and their effects on the accuracy of the global model have been discussed concerning untargeted poisoning attacks [Alkhunaizi et al., 2022].
@@ 4.4 Security
MI data are personal information that requires particular protection under specific confidentiality procedures. Any violation of the security procedure is a significant issue that might lead to the disclosure of personality [Truong et al., 2021]. The FL naturally provides a privacy protection mechanism by storing institutional datasets locally within each site without sharing them with third parties. However, the FL-based system can involve participants with unknown motives, and therefore, unduly entrusted clients can compromise security by reverse engineering [Li et al., 2020d, Darzidehkalani et al., 2022, Sohan and Basalamah, 2023]. Specifically, since a DL algorithm stores information in its weights, the untrusted client may reconstruct some portion of the data in a decentralized network [Darzidehkalani et al., 2022]. For example, recent studies have successfully rendered a face from MRI brain data [Schwarz et al., 2019, Abramian and Eklund, 2019].

### v2

@@ 2.2 Descriptive statistical analysis
The review integrates applications of FL with a wide range of MI modalities, particularly 63 papers related to MRI, 24 to CT, 23 to X-ray imaging, and 17 articles linked to histopathological imaging. In particular, 10 articles examined multimodal imaging techniques that involve CT and MRI, CT and PET, CT and X-ray, MRI and PET, ultrasonography and mammography, or CT, MRI, ultrasonography and histopathological images. The remaining articles covers PET, OCT, mammography, dermatoscopic, retinal, endoscopic and microscopy imaging. Fig. 4A. provides an exact distribution of articles involving specific topics. Fig. 4B shows a distribution of the papers examined according to the task they address. Among them, 68 articles focus on classification, 46 on segmentation, and 20 on detection. The rest of the papers broadly address prediction, prognosis, reconstruction, translation, restoration, normalization, denoising, screening, and super-resolution, with a few considering multiple tasks. Figure 5 summarizes the most commonly used aggregation and learning methods, categorized by imaging modality and deep learning task in the reviewed articles. Notably, the simplest and most established aggregation technique, Federated Averaging (FedAvg), is commonly employed across all modalities and deep learning tasks considered in this review. Considering the learning procedure, the traditional deep learning approaches remain predominant across modalities and medical imaging tasks. A detailed description of aggregation and learning techniques can be found in sections 3.3, 5 and 6. Additionally, Table 3 presents a comparative and detailed information gathered from the evaluated articles, including tasks, employed datasets with references, the presence of heterogeneous data, and the number of clients involved in the analysis.
@@
Cheng et al. [2025] Segmentation Dataset from article [Kuijf et al., 2019], own dataset Yes 7
Narmadha and Varalakshmi [2025] Segmentation Brain MRI segmentation, Brain Tumor Segmentation datasets N/A 5, 7, 10
@@
Li et al. [2025b] Classification Own dataset Yes 3
IXI dataset: https://brain-development.org/ixi-dataset/
Brain MRI segmentation dataset: https://www.kaggle.com/datasets/mateuszbuda/lgg-mri-segmentation
Brain Tumor Segmentation dataset: https://www.kaggle.com/datasets/nikhilroxtomar/brain-tumor-segmentation
Table 3: Summary of federated learning frameworks employed in various medical imaging tasks and imaging modalities. The column “Heterogeneous data” indicates whether heterogeneous data were used in the aggregation procedure. The last column refers to the number of clients involved in the deep learning task. All links were verified as of December 19, 2025.
Reference Task Used datasets Heterogeneous Number
data of clients
Weng et al. [2025] Classification ABIDE [Di Martino et al., 2014] N/A 4
Song et al. [2025] Classification ADNI [Jack Jr et al., 2008], AIBL [Ellis et al., 2009], own dataset Yes 3
Zhao et al. [2025] Classification REST-meta-MDD [Yan et al., 2019] Yes 4
Onaizah et al. [2025] Classification Dataset from Cheng et al. [2015] N/A 10, 20, 30, 50
Appasami and Savarimuthu [2025] Classification Own dataset N/A 4
@@
Wu et al. [2024] Reconstruction fastMRI [Knoll et al., 2020], CC359 [Beauferris et al., 2022], own dataset Yes 3
@@
Nezhad et al. [2025] Reconstruction fastMRI [Knoll et al., 2020], dataset from Elmas et al. [2023], Souza et al. [2018] Yes 3
Ahmed et al. [2025] Reconstruction fastMRI [Knoll et al., 2020], IXI Yes 3, 5, 10
@@
Wu et al. [2023] Classification RSNA ICH Detection [Flanders et al., 2020] Yes 20
@@
Yang et al. [2025] Denoising NIH-AAPM-Mayo Low-Dose CT [McCollough, 2016] N/A 8
@@
Zhou et al. [2023] Denoising Dataset from article [Xue et al., 2022], own dataset Yes 3
@@
COVID-19-9: https://figshare.com/authors/MedSeg/9940190
MS COVID-19-CT: https://sirm.org/COVID-19/
@@
LIDC-IDRI dataset: https://www.kaggle.com/datasets/zhangweiled/lidcidri
@@
Karmakar et al. [2025] Classification Lungs Disease Dataset (4 types) dataset, COVID-19 Radiography Database N/A 3, 5, 7
Lotfinia et al. [2025] Classification CheXpert [Irvin et al., 2019], PadChest [Bustos et al., 2020], ChestX-ray14 [Wang et al., 2017], VinDr-CXR [Nguyen et al., 2022b], PediCXR [Pham et al., 2023] Yes 5
@@
Dong and Voiculescu [2021] Detection CheXpert [Irvin et al., 2019], ChestX-ray8 [Wang et al., 2017], VinDr-CXR [Nguyen et al., 2022b] N/A 3, 6
@@
COVID-19 Radiography Database: https://www.kaggle.com/datasets/kumcs2004/covid-19-radiography-database
Table 6: Part IV.
Reference Task Used datasets Heterogeneous Number
data of clients
@@
Kumari et al. [2025] Classification BreakHis [Spanhol et al., 2015] Yes 5
Gupta et al. [2025] Classification BreakHis [Spanhol et al., 2015] Yes 5
@@
Wu et al. [2023] Classification HAM10000 [Tschandl et al., 2018], PH2 [Mendonça et al., 2013], Interactive atlas of dermoscopy, ISIC2019 [Codella et al., 2018, Tschandl et al., 2018, Combalia et al., 2019] Yes 5, 10
@@
Weng et al. [2025] Classification Derm7pt [Kawahara et al., 2018] N/A 3
@@
Dib [2025] Classification Diabetic Retinopathy Detection dataset Yes 2-10
Mahadi et al. [2025] Classification Eye Diseases Classification Dataset (EDC) [Wahab Sait, 2023] Yes 5
@@
Table 7: Part V.
Reference Task Used datasets Heterogeneous Number
data of clients
Gupta et al. [2025] Classification BUS [Al-Dhabyani et al., 2020] No 5
Hu et al. [2025] Classification Own dataset N/A 5
@@
Fiorentino et al. [2025] Detection Dataset from Burgos-Artizzu et al. [2020], Sendra-Balcells et al. [2023] 2 5
@@
Pervez et al. [2025] Classification Own dataset N/A 4
@@
Ma et al. [2025] Segmentation CT & MRI: CHAOS19 CT, MRI [Kavur et al., 2021] Yes 2
Raggio et al. [2025] Synthesis CT & MRI: Own dataset Yes 4
@@ 3.1 Theory
where $\mathcal{L}$ denotes the global loss function that is computed by combining $C$ local losses $\{\mathcal{L}_{c}\colon c\in\{1,2,...,C\}\}$, each contributing under normalized weight $w_{c}\geq 0$, and $\mathcal{D}_{c}=\left\{\mathbf{x}_{n},\mathbf{y}_{n}\right\}_{n=1}^{n_{c}}$ is the private dataset available only for cc–th client with training data $\mathbf{x}_{n}$ and corresponding labels $\mathbf{y}_{n}$.
@@ 3.3 Optimization and aggregation
Assuming a local (single) update for client cc–th performed during one local epoch is given by $\mathbf{\Theta}_{c}^{r}=\mathbf{\Theta}^{r}-\eta\nabla\mathcal{L}_{c}(\mathbf{\Theta}^{r};\mathcal{D}_{c})$, the Eq. (2) can be rewritten as:
@@
[…]
@@
where $\mathbf{\Theta}_{c,l}^{r}$ refers to the parameters at a global round $r$ from ll–th non-normalization layer in client $c$.
@@ 3.4 Learning
Vertical FL, also known as heterogeneous FL, handles different feature spaces from the data (e.g., various medical modalities) but with overlapping identifiers, meaning that the same patient appears in all sets [Mothukuri et al., 2021, Brecko et al., 2022, Abad et al., 2022]. It can facilitate training a common distributed model, such as a risk management model [Brecko et al., 2022]. Implementing vertical FL is more challenging than horizontal FL, as it typically involves three types of participants: active, passive, and coordinator. The active participant is the institution that wants to build the model and provide data with both the sample set and their labels. The passive participant only has the sample set, and its feature set differs from the active participant’s. The coordinator is a trusted third party responsible for coordinating training and communication among the participants. In some cases, the active participant is also the coordinator [Li et al., 2023b]. Furthermore, vertical FL requires encryption and a specific learning protocol, as the training process differs from horizontal FL [Mothukuri et al., 2021, Brecko et al., 2022]. Firstly, it is essential to align the training data in order to identify common sample identifiers. After alignment, the participants begin training the model. The most common training protocol is gradient descent, which involves transmitting intermediate values (e.g., softmax after the forward computation). Passive participants send the encoded intermediate results to the active participant. The active participant then calculates the losses and combines them with its loss. The aggregated loss is sent back to the passive participants. Active and passive participants use this aggregated loss to perform backpropagation and update the gradient for their local models. The encrypted gradients are sent to the coordinator, which decrypts them using a private key and returns them to the individual participants. The participants then retrain the model, and this entire process is iteratively repeated until the convergence of the model or a predefined condition is satisfied [Li et al., 2023b, Liu et al., 2024b]
@@ 4.1 Model accuracy
The bias that favors a specific client is a factor that could affect the accuracy of the global model [Darzidehkalani et al., 2022b]. If such a situation occurs, it can lead to an increase in the accuracy of the DL task for one client and a decrease in the accuracy for the others. The reason why the bias appears is the variability in the size and distribution of the datasets across the involved sites. The models generated by participants with larger datasets might be assigned with higher weighting factors, thus resulting in an over-representation or under-representation of specific patterns [Abay et al., 2020]. Besides, depending on the aggregation technique, only some clients can be involved, e.g., the institutions that conduct the entire local training within the specific timeframe. This can lead the FL algorithm to favor clients with a better computing infrastructure or faster and more reliable network connections. In clinical contexts, this usually corresponds to larger and richly endowed imaging centres, further reinforcing site-specific biases.
@@
4.2 Data heterogeneity
Data heterogeneity is a factor that significantly affects the accuracy of the global model in MI. Imaging data across institutions are typically collected with the equipment provided by various vendors under different acquisition protocols or even various data reconstruction procedures [Gibson et al., 2018, Jiang et al., 2022, Darzidehkalani et al., 2022b]. For example, histological images may look different due to various staining conditions. The MRI data may exhibit variations in distributional properties due to multifarious imaging protocols or contrasts applied, such as $T_{1}$-$T_{2}$-weighting or diffusion-weighting [Jiang et al., 2022], or even the procedure used to reconstruct magnitude data resulting in various statistical properties of the data [Den Dekker and Sijbers, 2014, Pieciak et al., 2016, 2017, Vegas-Sánchez-Ferrero et al., 2017]. Consequently, the data across the institutions may be nonindependent and nonidentically distributed (non-IID).
@@ 4.2 Data heterogeneity
As mentioned above, data heterogeneity originates from the absence of unified imaging standardization procedures. Although the involved sites use the same acquisition protocol to collect data, they may employ different software to preprocess data, leading, for example, to reduced reproducibility of experiments [Veraart et al., 2023]. In addition to these, the number of samples and their distribution can vary over time [Darzidehkalani et al., 2022b, Aouedi et al., 2022]. For example, the number of patients affected by a particular disease in one hospital can fluctuate, including the patient’s recovery, death, or transfer of care. Finally, the institutional data deviates according to the demography of the sample. As an illustration, the MRI data change to such an extent throughout life that it allows reconstruction of the patient’s age [Cox et al., 2016, Peng et al., 2021, Pieciak et al., 2023]. Finally, the diverse nature of the data can result in reduced global model accuracy due to its inability to converge. The global model may perform exceptionally well for some clients but inaccurately for others [Darzidehkalani et al., 2022b]. According to Zhao et al. [2018], highly heterogeneous data can cause up to 50% degradation in the accuracy of the FL global model, making heterogeneous data one of the most critical issues to address. In the context of MI, such a severe decrease in performance quality has not been observed. Nevertheless, the authors reported a decline of approximately five percentage points in classification accuracy [Zhao and Huang, 2023], a reduction in the Sørensen–Dice index ranging from about 0.05 to 0.15 for segmentation tasks [Jiang et al., 2023, Yang et al., 2023], and a decrease in the MSSIM index by approximately 0.17 for data reconstruction task [Feng et al., 2023].
In addition to this, institutions still lack proper infrastructure for image data processing and data management systems that are compliant with current standards [Darzidehkalani et al., 2022b].
@@ 4.3 Malicious clients
The quality deterioration of the results generated from the global model may come from malicious clients in all institutions. Due to its distributed nature, the FL is susceptible to model poisoning attacks. These attacks occur if a malicious client sends manipulated model updates to the server [Alkhunaizi et al., 2022, Zhang et al., 2022c, Onsu et al., 2023]. Model poisoning attacks can be classified into two categories: untargeted and targeted. In the former case, a corrupted global model generates incorrect predictions for multiple test inputs. In the latter scenario, a corrupted global model causes incorrect predictions for specific test inputs chosen by the attacker, while the accuracy of the global model for other inputs generally remains unaffected [Abad et al., 2022, Zhang et al., 2022c]. In clinical settings, the behavior of malicious clients may be unintentional rather than adversarial. It can occur due to issues such as corrupted imaging data, misconfigured preprocessing pipelines, or systematic annotation errors at individual sites. Despite being unintentional, these issues can have a disproportionately negative impact on the aggregated model. For MI applications, the risks posed by malicious clients and their effects on the accuracy of the global model have been discussed concerning untargeted poisoning attacks [Alkhunaizi et al., 2022].
@@ 4.4 Security
MI data are personal information that requires particular protection under specific confidentiality procedures. Any violation of the security procedure is a significant issue that might lead to the disclosure of personality [Truong et al., 2021]. The FL naturally provides a privacy protection mechanism by storing institutional datasets locally within each site without sharing them with third parties. However, the FL-based system can involve participants with unknown motives, and therefore, unduly entrusted clients can compromise security by reverse engineering [Li et al., 2020d, Darzidehkalani et al., 2022b, Sohan and Basalamah, 2023]. Specifically, since a DL algorithm stores information in its weights, the untrusted client may reconstruct some portion of the data in a decentralized network [Darzidehkalani et al., 2022b]. For example, recent studies have successfully rendered a face from MRI brain data [Schwarz et al., 2019, Abramian and Eklund, 2019].